# Disambiguation in Conversational Question Answering in the Era of LLM: A Survey

### v1

@@ Abstract
Ambiguity remains a fundamental challenge in Natural Language Processing (NLP) due to the inherent complexity and flexibility of human language. With the advent of Large Language Models (LLMs), addressing ambiguity has become even more critical due to their expanded capabilities and applications. In the context of Conversational Question Answering (CQA), this paper explores the definition, forms, and implications of ambiguity for language driven systems, particularly in the context of LLMs. We define key terms and concepts, categorize various disambiguation approaches enabled by LLMs, and provide a comparative analysis of their advantages and disadvantages. We also explore publicly available datasets for benchmarking ambiguity detection and resolution techniques and highlight their relevance for ongoing research. Finally, we identify open problems and future research directions, proposing areas for further investigation. By offering a comprehensive review of current research on ambiguities and disambiguation with LLMs, we aim to contribute to the development of more robust and reliable language systems.
Disambiguation in Conversational Question Answering in the Era of LLM: A Survey
@@ 1 Introduction
Organization of this Survey. We structure this survey around three core research questions (see Figure 1): RQ1: Why do ambiguities arise in language, and how can we detect them? RQ2: How can we disambiguate, particularly using LLMs? RQ3: How can we automate disambiguation strategies in real-world applications? Section 2 addresses RQ1 by defining key concepts, presenting a taxonomy, and reviewing ambiguity detection methods. Section 3 tackles RQ2 by categorizing LLM-based disambiguation approaches and analyzing their strengths and weaknesses. To support these, Section 4 surveys relevant public datasets used for benchmarking. Finally, Section 5 explores open challenges and outlines future directions, centering on RQ3: how to orchestrate disambiguation effectively in practice.
@@ 2.2 Taxonomy of Ambiguity
In Table 1, we present a comparative analysis of these taxonomies to highlight common grounds despite their differences. To cater to broader applications and provide clarity, we propose simplifying existing taxonomies into three overarching categories. We argue that these categories can encompass all existing taxonomies, irrespective of the underlying tasks, thereby offering a more cohesive framework for understanding ambiguities.
@@
Semantic Ambiguity: When a sentence is grammatically correct but semantically unclear, due to ambiguity in a word, phrase, or the overall interpretation. The more common case involves ambiguity at the word or phrase level, often referred to as lexical ambiguity (Navigli, 2009; Beekhuizen et al., 2021), where a term has multiple possible meanings. As shown in Table 1, this type is listed across most prior work, with the exception of Zhang and Choi (2023), where they mention it as ‘word sense disambiguation.’ Similarly, the ‘Unfamiliar’ category in Zhang et al. (2024c) aligns with this type, as unknown words are inherently open to interpretation until contextual or domain-specific knowledge is applied. Beyond word-level issues, semantic ambiguity can also stem from interpretive variation at the sentence level. This includes the usage of literal vs. pragmatic words as mentioned by Liu et al. (2023), who refer to it as pragmatic ambiguity, and ‘Literal vs. Implied Interpretations’ by Zhang and Choi (2023). The ‘Figurative’ type in Liu et al. (2023) also falls into this category, as does the ‘Contradiction’ category in Zhang et al. (2024c) because of conflicts with the semantics of previous statements.
@@ 2.3 Ambiguity Detection
The body of work for detecting ambiguity can be broadly categorized into three major groups: traditional methods (not language model-based), language model-based methods, and large language model (LLM)-based methods. In Table 2, we summarize each method’s approach, model inputs, and the types of ambiguity it addresses based on our taxonomy. We give more details below.
@@
Despite the flexibility of LLMs, these works collectively show that ambiguity detection—particularly fine-grained distinctions among types—remains a complex problem. We will revisit these challenges in Section 5.
Figure 2: We find existing disambiguation efforts using LLMs broadly fall into these three major categories: Left. Query Rewriting, Middle. Long Form Answer Generation, Right. Asking Clarifying Questions. These policies have different requirements and also work in different processing steps in CQA pipeline, resulting in unique advantages and disadvantages for each approach. We highlight a couple here and provide a more comprehensive list in Table 5.
@@ 3 How To Disambiguate?
In the era of LLMs, disambiguation is gaining increasing attention due to their extensive world knowledge and advanced capabilities, surpassing traditional and smaller language models. However, current research in this area often lacks systematic categorization and tends to address various aspects in isolation. To that end, in this paper, we argue existing disambiguation works fall in three major policies, which we present in Figure 2. We describe each of them below.
@@ 3.1 Query Rewriting (QR)
The emergence of LLMs has enabled more advanced query reformulation, moving beyond term-based edits to deeper semantic understanding and contextual refinement across downstream tasks Wang et al. (2023). Recent research works, such as Ma et al. (2023); Jagerman et al. (2023), have demonstrated the efficacy of LLM-based query reformulation in zero-shot and few-shot settings, particularly valuable when domain-specific training data is scarce. The principled way of QR is shown in Figure 2 (Left), where an LLM is prompted with previous chat history and other relevant information as context. Some advanced prompting, such as Ye et al. (2023) also includes “rewrite-then-edit” framework. Apart from prompting, LLMs also have been fined-tuned Peng et al. (2024) or used to generate Supervised Fine Tuning (SFT) dataset to improve QR model either through a re-ranker Mao et al. (2024) or preference optimization Zhang et al. (2024b).
@@ 3.2 Long Form Answer Generation (LFAG)
Generating long-form answers to ambiguous questions involves presenting all valid interpretations alongside their corresponding answers. For instance, the question “Who has the highest goals in world football?” can refer to either men’s or women’s football. A well-structured response would be: “Ali Daei holds the record in men’s football, while Christine Sinclair does in women’s football.” As shown in Figure 2 (Middle), this task typically comprises three steps: 1) Disambiguating the question, 2) Answering each interpretation, and 3) Consolidating the results into a single, coherent response. Early methods streamlined these steps into a single model inference. Stelmakh et al. (2022) finetuned T5 to directly produce long-form answers. More recent LLM-based approaches, such as Gao et al. (2023), show that few-shot prompting can be similarly effective without fine-tuning. To reduce reasoning load, Amplayo et al. (2022) proposed a two-step method: first inferring multiple interpretations, then generating a long-form answer from them. RAC Kim et al. (2023) introduced retrieval-augmented disambiguation to generate answers with supporting evidence (Steps 1–2), while ToC Kim et al. (2023) extended this via iterative retrieval to capture overlooked interpretations, trading off efficiency. DIVA In et al. (2024) improved efficiency by modeling a reasoning chain that compresses this process into a single step, maintaining performance while reducing complexity.
@@ 3.3 Asking Clarifying Question (ACQ)
This is one of the most extensively studied disambiguation policies, with approaches ranging from rule-based prompts (e.g., “Did you mean A or B?” Coden et al. (2015), “What do you want to know about QUERY?” Zamani et al. (2020), or category-based options Lee et al. (2023)) to traditional machine learning Zhang et al. (2018); Rao and Daumé III (2018, 2019) and language model-based methods Xu et al. (2019); Aliannejadi et al. (2019). Several works also introduce new datasets Xu et al. (2019); Kumar and black (2020); Min et al. (2020); Guo et al. (2021), discussed further in Section 4. However, these methods often struggle with complex queries and rely on annotated corpora, which could be difficult to obtain.
With the advent of LLMs, recent studies have leveraged prompt-based approaches Kuhn et al. (2022); Deng et al. (2023b); Zhang et al. (2024c), typically employing zero-shot or few-shot Chain-of-Thought (CoT) prompting strategies. These methods mirror the Long-form Answer Generation pipeline but focus on analyzing multiple valid interpretations to generate clarifying questions, as shown in Figure 2 (Right). Like QR, they reduce the need for domain-specific data and can be training-free while supporting complex question structures. Some works adopt a two-stage pipeline: first detecting ambiguity, then generating suitable clarification questions. For instance, Zhang and Choi (2023) proposed an innovative uncertainty estimation technique for ambiguity detection that quantifies intent entropy through simulated user-assistant interactions. Finally, similar to QR, LLMs can be also be fine-tuned to generate clarifying questions. For example, Zhang et al. (2024a); Kim et al. (2024) fine-tuned various LLMs, such as Llama-2-7B Touvron et al. (2023), Gemma-7B Team et al. (2024), and Llama-3-8B Dubey et al. (2024).
Table 3 summarizes how disambiguation techniques address different ambiguity types. QR handles all three by reformulating queries to fix syntactic issues, resolve semantic confusion through inferred interpretations, and incorporate missing contextual details from prior conversation. LFAG handles semantic and contextual ambiguity by presenting multiple plausible interpretations, including those that differ semantically as well as those that are plausible when considering different contexts. ACQ resolves semantic and contextual ambiguity by explicitly asking the user to confirm among similar options or supply missing information. While QR might look most appealing for its broad coverage, it still faces key challenges such as semantic drift Anand et al. (2023) and practical concerns like latency, cost, and error propagation in production Tanjim et al. (2025). We will discuss the strengths and limitations of each approach further in Section 5.
@@ 4 Benchmarks
Ambiguity Detection and ACQ. Most existing datasets related to ambiguity fall into the category of detecting the need for clarification and necessary disambiguation by asking clarification questions. Notable datasets in this area include CLAQUA Xu et al. (2019), ClarQ Kumar and black (2020), AmbigNQ Min et al. (2020), ClariQ Aliannejadi et al. (2020), Abg-CoQA Guo et al. (2021), PACIFIC Deng et al. (2022), CAmbigNQ Lee et al. (2023), and CLAMBER Zhang et al. (2024c). These corpora exhibit significant variation across several dimensions, each contributing uniquely to the understanding of ambiguities in dialogue systems, as listed in Table 4. Among them, the CLAMBER benchmark Zhang et al. (2024c) has emerged as the first comprehensive evaluation benchmark for LLM-based ambiguity detection and ACQ, providing valuable insights into the current limitations of LLM-based approaches and establishing baseline metrics for future research. Statistics for all these datasets, along with their corresponding URLs, appear in Table 4. Metrics typically used for ambiguity detection include classification metrics such as Precision, Recall, F1, Accuracy, and AUROC score Zhang et al. (2024c); Tanjim et al. (2025). For ACQ, the metrics are usually automatic text evaluation metrics, such as BLEU Papineni et al. (2002) or ROUGE Lin (2004). However, some studies criticize the limitations of these metrics and favor human judgment instead Zamani et al. (2020).
Query Rewriting. There are two prominent benchmark datasets for evaluating the quality of rewritten queries. The pioneering dataset in this area is CANARD Elgohary et al. (2019), which includes questions with context and their rewritten versions. This was followed by QReCC Anantha et al. (2021), where each user question is accompanied by a human-rewritten query, and answers to questions within the same conversation may be distributed across multiple web pages. Notably, QReCC is used in recent LLM-based QR approaches such as Ye et al. (2023) and Zhang et al. (2024b). Both of these datasets, along with their statistics and URLs, are listed in Table 4. It is important to note that, unlike datasets related to ACQ, these datasets do not contain specific fields or labels explicitly indicating ‘ambiguity’ in queries. As for metrics, similar to ACQ, BLEU and ROUGE are popular choices for measuring the quality of rewritten queries. Additionally, since QR is often employed for IR tasks, standard IR metrics such as mean reciprocal rank (MRR), mean average precision (MAP), and Recall@k and Precision@k are used to evaluate whether the rewritten query retrieves the correct information Ma et al. (2023); Ye et al. (2023). For these purposes, popular open-domain QA datasets like NQ Kwiatkowski et al. (2019), TriviaQA Joshi et al. (2017), and HotpotQA Yang et al. (2018) are often used as benchmarks. However, we do not list them here as they do not focus specifically on ambiguity and lack corresponding human-rewritten queries.
@@
Long Form Answer Generation. To the best of our knowledge, ASQA Stelmakh et al. (2022) is the only dataset that falls into this category. ASQA is a long-form QA dataset derived from a subset of ambiguous questions in the AmbigNQ dataset Min et al. (2020). Its statistics and corresponding URL are provided in Table 4. The dataset is designed to evaluate how well systems can generate comprehensive answers that cover all valid interpretations. Two main metrics are used to assess the generation quality In et al. (2024): Disambig-F1 (D-F1) Stelmakh et al. (2022), which assesses the accuracy of responses by verifying correct answers to disambiguated questions using an F1 score, and ROUGE, which evaluates the correctness by comparing them to ground-truth long-form answers.
@@ 5 Open Problems and Challenges
Detecting Ambiguities. While LLMs have exceptional generative capabilities, recent studies consistently highlight the challenges of using LLMs to detect ambiguous queries with high performance. For example, Zhang and Choi (2023) achieved an AUROC of $0.570.57$ on AmbigNQ Min et al. (2020) using LLaMA-2-13B-Chat, while Zhang et al. (2024c) reported a best F1 score of $0.530.53$ on their dataset using GPT-3.5-Turbo. Tanjim et al. (2025) shares a similar study and highlight a relatively lower performance using GPT-3.5-Turbo and LLaMA-3.1-70B. One potential reason, as suggested by Liu et al. (2023), is that LLMs are not inherently designed to model ambiguities.

### v2

# Disambiguation in Conversational Question Answering in the Era of LLMs and Agents: A Survey

@@ Abstract
Ambiguity remains a fundamental challenge in Natural Language Processing (NLP) due to the inherent complexity and flexibility of human language. With the advent of Large Language Models (LLMs), addressing ambiguity has become even more critical due to their expanded capabilities and applications. In the context of Conversational Question Answering (CQA), this paper explores the definition, forms, and implications of ambiguity for language driven systems, particularly in the context of LLMs. We define key terms and concepts, categorize various disambiguation approaches enabled by LLMs, and provide a comparative analysis of their advantages and disadvantages. We also explore publicly available datasets for benchmarking ambiguity detection and resolution techniques and highlight their relevance for ongoing research. Finally, we identify open problems and future research directions, especially in agentic settings, proposing areas for further investigation. By offering a comprehensive review of current research on ambiguities and disambiguation with LLMs, we aim to contribute to the development of more robust and reliable LLM-based systems.
Disambiguation in Conversational Question Answering in the Era of LLMs and Agents: A Survey
@@ 1 Introduction
Organization of this Survey. We structure this survey around three core research questions (see Figure˜1): RQ1: Why do ambiguities arise in language, and how can we detect them? RQ2: How can we disambiguate, particularly using LLMs? RQ3: How can we automate disambiguation strategies in real-world applications? Section 2 addresses RQ1 by defining key concepts, presenting a taxonomy, and reviewing ambiguity detection methods. Section 3 tackles RQ2 by categorizing LLM-based disambiguation approaches and analyzing their strengths and weaknesses. To support these, Section 4 surveys relevant public datasets used for benchmarking. Finally, Section 5 explores open challenges and outlines future directions, centering on RQ3: how to orchestrate disambiguation effectively in practice.
@@ 2.2 Taxonomy of Ambiguity
In Table˜1, we present a comparative analysis of these taxonomies to highlight common grounds despite their differences. To cater to broader applications and provide clarity, we propose simplifying existing taxonomies into three overarching categories. We argue that these categories can encompass all existing taxonomies, irrespective of the underlying tasks, thereby offering a more cohesive framework for understanding ambiguities.
@@
Semantic Ambiguity: When a sentence is grammatically correct but semantically unclear, due to ambiguity in a word, phrase, or the overall interpretation. The more common case involves ambiguity at the word or phrase level, often referred to as lexical ambiguity (Navigli, 2009; Beekhuizen et al., 2021), where a term has multiple possible meanings. As shown in Table˜1, this type is listed across most prior work, with the exception of Zhang and Choi (2023), where they mention it as ‘word sense disambiguation.’ Similarly, the ‘Unfamiliar’ category in Zhang et al. (2024c) aligns with this type, as unknown words are inherently open to interpretation until contextual or domain-specific knowledge is applied. Beyond word-level issues, semantic ambiguity can also stem from interpretive variation at the sentence level. This includes the usage of literal vs. pragmatic words as mentioned by Liu et al. (2023), who refer to it as pragmatic ambiguity, and ‘Literal vs. Implied Interpretations’ by Zhang and Choi (2023). The ‘Figurative’ type in Liu et al. (2023) also falls into this category, as does the ‘Contradiction’ category in Zhang et al. (2024c) because of conflicts with the semantics of previous statements.
@@ 2.3 Ambiguity Detection
The body of work for detecting ambiguity can be broadly categorized into three major groups: traditional methods (not language model-based), language model-based methods, and large language model (LLM)-based methods. In Table˜2, we summarize each method’s approach, model inputs, and the types of ambiguity it addresses based on our taxonomy. We give more details below.
@@
Despite the flexibility of LLMs, these works collectively show that ambiguity detection—particularly fine-grained distinctions among types—remains a complex problem. We will revisit these challenges in Section˜5.
Figure 2: We find existing disambiguation efforts using LLMs broadly fall into these three major categories: Left. Query Rewriting, Middle. Long Form Answer Generation, Right. Asking Clarifying Questions. These policies have different requirements and also work in different processing steps in CQA pipeline, resulting in unique advantages and disadvantages for each approach. We highlight a couple here and provide a more comprehensive list in Table˜5.
@@ 3 How To Disambiguate?
In the era of LLMs, disambiguation is gaining increasing attention due to their extensive world knowledge and advanced capabilities, surpassing traditional and smaller language models. However, current research in this area often lacks systematic categorization and tends to address various aspects in isolation. To that end, in this paper, we argue existing disambiguation works fall in three major policies, which we present in Figure˜2. We describe each of them below.
@@ 3.1 Query Rewriting (QR)
The emergence of LLMs has enabled more advanced query reformulation, moving beyond term-based edits to deeper semantic understanding and contextual refinement across downstream tasks Wang et al. (2023). Recent research works, such as Ma et al. (2023); Jagerman et al. (2023), have demonstrated the efficacy of LLM-based query reformulation in zero-shot and few-shot settings, particularly valuable when domain-specific training data is scarce. The principled way of QR is shown in Figure˜2 (Left), where an LLM is prompted with previous chat history and other relevant information as context. Some advanced prompting, such as Ye et al. (2023) also includes “rewrite-then-edit” framework. Apart from prompting, LLMs also have been fined-tuned Peng et al. (2024) or used to generate Supervised Fine Tuning (SFT) dataset to improve QR model either through a re-ranker Mao et al. (2024) or preference optimization Zhang et al. (2024b).
@@ 3.2 Long Form Answer Generation (LFAG)
Generating long-form answers to ambiguous questions involves presenting all valid interpretations alongside their corresponding answers. For instance, the question “Who has the highest goals in world football?” can refer to either men’s or women’s football. A well-structured response would be: “Ali Daei holds the record in men’s football, while Christine Sinclair does in women’s football.” As shown in Figure˜2 (Middle), this task typically comprises three steps: 1) Disambiguating the question, 2) Answering each interpretation, and 3) Consolidating the results into a single, coherent response. Early methods streamlined these steps into a single model inference. Stelmakh et al. (2022) finetuned T5 to directly produce long-form answers. More recent LLM-based approaches, such as Gao et al. (2023), show that few-shot prompting can be similarly effective without fine-tuning. To reduce reasoning load, Amplayo et al. (2022) proposed a two-step method: first inferring multiple interpretations, then generating a long-form answer from them. RAC Kim et al. (2023a) introduced retrieval-augmented disambiguation to generate answers with supporting evidence (Steps 1–2), while ToC Kim et al. (2023a) extended this via iterative retrieval to capture overlooked interpretations, trading off efficiency. DIVA In et al. (2024) improved efficiency by modeling a reasoning chain that compresses this process into a single step, maintaining performance while reducing complexity.
@@ 3.3 Asking Clarifying Question (ACQ)
This is one of the most extensively studied disambiguation policies, with approaches ranging from rule-based prompts (e.g., “Did you mean A or B?” Coden et al. (2015), “What do you want to know about QUERY?” Zamani et al. (2020), or category-based options Lee et al. (2023)) to traditional machine learning Zhang et al. (2018); Rao and Daumé III (2018, 2019) and language model-based methods Xu et al. (2019); Aliannejadi et al. (2019). Several works also introduce new datasets Xu et al. (2019); Kumar and black (2020); Min et al. (2020); Guo et al. (2021), discussed further in Section˜4. However, these methods often struggle with complex queries and rely on annotated corpora, which could be difficult to obtain.
With the advent of LLMs, recent studies have leveraged prompt-based approaches Kuhn et al. (2022); Deng et al. (2023b); Zhang et al. (2024c), typically employing zero-shot or few-shot Chain-of-Thought (CoT) prompting strategies. These methods mirror the Long-form Answer Generation pipeline but focus on analyzing multiple valid interpretations to generate clarifying questions, as shown in Figure˜2 (Right). Like QR, they reduce the need for domain-specific data and can be training-free while supporting complex question structures. Some works adopt a two-stage pipeline: first detecting ambiguity, then generating suitable clarification questions. For instance, Zhang and Choi (2023) proposed an innovative uncertainty estimation technique for ambiguity detection that quantifies intent entropy through simulated user-assistant interactions. Finally, similar to QR, LLMs can be also be fine-tuned to generate clarifying questions. For example, Zhang et al. (2024a); Kim et al. (2024) fine-tuned various LLMs, such as Llama-2-7B Touvron et al. (2023), Gemma-7B Team et al. (2024), and Llama-3-8B Dubey et al. (2024).
Table˜3 summarizes how disambiguation techniques address different ambiguity types. QR handles all three by reformulating queries to fix syntactic issues, resolve semantic confusion through inferred interpretations, and incorporate missing contextual details from prior conversation. LFAG handles semantic and contextual ambiguity by presenting multiple plausible interpretations, including those that differ semantically as well as those that are plausible when considering different contexts. ACQ resolves semantic and contextual ambiguity by explicitly asking the user to confirm among similar options or supply missing information. While QR might look most appealing for its broad coverage, it still faces key challenges such as semantic drift Anand et al. (2023) and practical concerns like latency, cost, and error propagation in production Tanjim et al. (2025). We will discuss the strengths and limitations of each approach further in Section˜5.
@@ 4 Benchmarks
Ambiguity Detection and ACQ. Most existing datasets related to ambiguity fall into the category of detecting the need for clarification and necessary disambiguation by asking clarification questions. Notable datasets in this area include CLAQUA Xu et al. (2019), ClarQ Kumar and black (2020), AmbigNQ Min et al. (2020), ClariQ Aliannejadi et al. (2020), Abg-CoQA Guo et al. (2021), PACIFIC Deng et al. (2022), CAmbigNQ Lee et al. (2023), and CLAMBER Zhang et al. (2024c). These corpora exhibit significant variation across several dimensions, each contributing uniquely to the understanding of ambiguities in dialogue systems, as listed in Table˜4. Among them, the CLAMBER benchmark Zhang et al. (2024c) has emerged as the first comprehensive evaluation benchmark for LLM-based ambiguity detection and ACQ, providing valuable insights into the current limitations of LLM-based approaches and establishing baseline metrics for future research. Statistics for all these datasets, along with their corresponding URLs, appear in Table˜4. Metrics typically used for ambiguity detection include classification metrics such as Precision, Recall, F1, Accuracy, and AUROC score Zhang et al. (2024c); Tanjim et al. (2025). For ACQ, the metrics are usually automatic text evaluation metrics, such as BLEU Papineni et al. (2002) or ROUGE Lin (2004). However, some studies criticize the limitations of these metrics and favor human judgment instead Zamani et al. (2020).
Query Rewriting. There are two prominent benchmark datasets for evaluating the quality of rewritten queries. The pioneering dataset in this area is CANARD Elgohary et al. (2019), which includes questions with context and their rewritten versions. This was followed by QReCC Anantha et al. (2021), where each user question is accompanied by a human-rewritten query, and answers to questions within the same conversation may be distributed across multiple web pages. Notably, QReCC is used in recent LLM-based QR approaches such as Ye et al. (2023) and Zhang et al. (2024b). Both of these datasets, along with their statistics and URLs, are listed in Table˜4. It is important to note that, unlike datasets related to ACQ, these datasets do not contain specific fields or labels explicitly indicating ‘ambiguity’ in queries. As for metrics, similar to ACQ, BLEU and ROUGE are popular choices for measuring the quality of rewritten queries. Additionally, since QR is often employed for IR tasks, standard IR metrics such as mean reciprocal rank (MRR), mean average precision (MAP), and Recall@k and Precision@k are used to evaluate whether the rewritten query retrieves the correct information Ma et al. (2023); Ye et al. (2023). For these purposes, popular open-domain QA datasets like NQ Kwiatkowski et al. (2019), TriviaQA Joshi et al. (2017), and HotpotQA Yang et al. (2018) are often used as benchmarks. However, we do not list them here as they do not focus specifically on ambiguity and lack corresponding human-rewritten queries.
@@
Long Form Answer Generation. To the best of our knowledge, ASQA Stelmakh et al. (2022) is the only dataset that falls into this category. ASQA is a long-form QA dataset derived from a subset of ambiguous questions in the AmbigNQ dataset Min et al. (2020). Its statistics and corresponding URL are provided in Table˜4. The dataset is designed to evaluate how well systems can generate comprehensive answers that cover all valid interpretations. Two main metrics are used to assess the generation quality In et al. (2024): Disambig-F1 (D-F1) Stelmakh et al. (2022), which assesses the accuracy of responses by verifying correct answers to disambiguated questions using an F1 score, and ROUGE, which evaluates the correctness by comparing them to ground-truth long-form answers.
@@ 5.1 Detecting Ambiguities
While LLMs have exceptional generative capabilities, recent studies consistently highlight the challenges of using LLMs to detect ambiguous queries with high performance. For example, Zhang and Choi (2023) achieved an AUROC of $0.57$ on AmbigNQ Min et al. (2020) using LLaMA-2-13B-Chat, while Zhang et al. (2024c) reported a best F1 score of $0.53$ on their dataset using GPT-3.5-Turbo. Tanjim et al. (2025) shares a similar study and highlight a relatively lower performance using GPT-3.5-Turbo and LLaMA-3.1-70B. One potential reason, as suggested by Liu et al. (2023), is that LLMs are not inherently designed to model ambiguities.
5.2 How To Orchestrate?